For the eigenstate with n=o,1, and 2 ,compute the probability that the coordinate of a linear harmonic oscillator in its ground state has value greater than oscillator than the amplitude of a classical oscillator of the same energy level
1
Expert's answer
2020-05-11T20:04:14-0400

In quantum case "E_n=\Big(n+\frac{1}{2}\Big)\hbar\omega"

In classic case "E=kx^2/2=m\omega^2A^2/2"

Let's find the amplitude "A" of oscillations for a classical oscillator with energy equal to the energy of a quantum oscillator in the quantum state "n".

We obtain "E_n=m\omega^2A^2_n/2 \rightsquigarrow A_n=\sqrt{(2n+1)\frac{\hbar}{m\omega}}"


In quantum case

"\psi_n(x)=\frac{1}{\sqrt{2^n n!}}\cdot (\frac{m\omega }{\pi \hbar })^{1/4}\cdot e^{-\frac{m\omega x^2}{2\hbar}}\cdot H_n(\sqrt{\frac{m\omega }{\hbar }} x)"

where "H_n(y)" are Hermite polynomials, "H_0(y)=1" .


We want to compute the probability that the coordinate of a linear harmonic oscillator in its ground state (i.e. "n=0") has value greater than ''classic zone'' "x\in(-A,A)", i.e.


"\int\limits_{A_0}^\infty |\psi_0(x)|^2\,dx=\sqrt{\frac{m\omega}{\pi \hbar}}\int\limits_{A_0}^\infty e^{-m\omega x^2/\hbar}\,dx =""=\sqrt{\frac{m\omega}{\pi \hbar}} \cdot\frac{1}{2}\sqrt{\frac{\pi \hbar}{m \omega}}\operatorname{erfc}(1) \approx 0.0787"

Or approximately "7.87 \%"